Following is the profit function for a firm selling two products:
π=50X-X2+100Y-4Y2
Where X and Y represent output rates of products 1 and 2, respectively
Determine the profit maximizing output rates for the two products and the profit associated with these output rates
1
Expert's answer
2020-06-16T12:15:35-0400
"\frac {\delta {\pi}}{\delta x}=50-2x"


"\frac {\delta {\pi}}{\delta x}=100-8y"


"50-2x=0"


"x=25"


"100-8y=0"


"y=12.5"